solve for x; if


log 7(x3+7)-log7(x+3)=2

Grade:12

1 Answers

vikas askiitian expert
13 years ago

log 7(x3+7)-log7(x+3)=2

log(x3+7)/log7 - logx+3/log7 = 2

logx3+7 - logx+3 = 2log7

logx3+7/x+3 = log49

x3+7 = 49x + 49*3

x3 -49x - 20*7 = 0

x3 - 49x - 140 = 0

now solve for x